OpenStudy (anonymous):

A man has 18 coins in his pocket, all of which are dimes and quarters. If the total value of his change is 360 cents, how many dimes and how many quarters does he have? Your answer is number of dimes equals________ number of quarters equals______

OpenStudy (danjs):

12 coins = quarters + dimes 360 cents = 25q + 10d

OpenStudy (danjs):

2 eq / 2 unk

OpenStudy (danjs):

18 coins in the first , i'm sorry

OpenStudy (danjs):

18 = q + d 360 = 25q + 10d

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

25x+10y=360 25x+10y=360 x+y=18 +-10x-10y=-180 y=Dime 15x=180 x=Quarter x=12 Now plug in x to get y.